OpenStudy (anonymous):

The submarine dove 75ft, rose 25ft, then dove 140ft, and finally ascended 85ft, What was the next change in depth?

OpenStudy (anonymous):

Strange question... Depths in order: 75 50 190 105

OpenStudy (anonymous):

what?

hero (hero):

75 - 25 = 50 50 + 140 = 190 190 - 85 = 105 I don't know what they mean by next change in depth though.

OpenStudy (cwrw238):

very odd